# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

### v1

@@ III-A Overall Framework
1:Input: historical data $\mathcal{D}$, dataset split factor $\eta\in(0,1)$, sliding window factor $\epsilon>=0$,
Input: data augmentation threshold $N$ and factor $\rho$
@@
6: /************** Data Augmentation (Optional) **************/
@@
As illustrated in Fig. 1, EvoGO comprises three key phases: data preparation, model training, and population generation. To mitigate the risk of overfitting caused by limited population data, the data preparation phase incorporates a pairwise data augmentation strategy, enhancing training diversity without introducing additional evaluation costs. In the model training phase, a composite model architecture and a tailored loss function are employed to align the generative training objectives with the goals of evolutionary optimization, thereby addressing the objective misalignment issue. Moreover, to eliminate reliance on handcrafted rules, EvoGO adopts an end-to-end generative paradigm, replacing the heuristic components commonly used in traditional GMEAs. Within this data-driven framework, data preparation replaces conventional environmental selection; model training substitutes for basic surrogate modeling; and population generation takes the place of heuristic-based reproduction.
@@ III-B Data Preparation
The data preparation phase in EvoGO aims to construct high-quality and structured training data for subsequent model training. Specifically, it (i) selects elite samples from historical populations to ensure surrogate accuracy, (ii) optionally augments the data to improve diversity under limited evaluations, and (iii) generates a paired dataset to provide directional learning signals for the generative model.
@@
  • Elite Selection: A high-quality dataset $\mathcal{D}_{\text{SM}}$ is constructed by selecting individuals with the highest fitness values from all historical populations $\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}$. This ensures accurate surrogate modeling by concentrating on representative and promising solutions.
  • Data Augmentation (Optional): To mitigate the limited data availability in expensive optimization scenarios, an optional learning-based data augmentation strategy is employed. Specially, a Variational AutoEncoder (VAE) is used to generate diverse synthetic samples that resemble the current population, thereby expanding the generative model dataset $\mathcal{D}_{\text{GM}}$ when the original population is insufficient.
  • Paired Set Construction: Following augmentation (if applied), the surrogate model estimates the fitness of all candidate solutions. The samples are then ranked and partitioned into superior ($\mathcal{D}_{\text{GM}^{+}}$) and inferior ($\mathcal{D}_{\text{GM}^{-}}$) subsets based on a predefined split ratio $\eta$. Their Cartesian product forms the paired training set $\mathcal{D}_{\text{GM}}$, enabling the composite model to learn both effective descent directions and adaptive step sizes.
@@ III-C Model Training
In the model training phase, we develop a composite model architecture that integrates a surrogate model with a pair of generative models. To support this structure, we define a dedicated composite loss function that enables directional learning and optimization-aware training. This phase is designed as an online process, where lightweight training is conducted during each evolutionary generation. Although the training involves three model components, only the generative model $\bm{\gamma}$ is used during population generation. The overall training procedure is described below, followed by detailed explanations of the model architecture and loss formulation.
@@
Conventional GMEAs typically rely on generative models in conjunction with heuristic procedures, often neglecting the optimization guidance available from surrogate-estimated objective landscapes. To address these limitations, we propose a composite model architecture (see Fig. 2) that integrates a surrogate model $\tilde{f}$ with a forward generative model $\bm{\gamma}$, which maps inferior solutions to superior ones, and a pseudo-inverse model $\bm{\gamma}^{\prime}$ that performs the reverse mapping. In contrast to GANs [74] and CycleGANs [75], which rely on discriminators, our approach leverages the surrogate model to impose optimization-aware learning via gradient back-propagation.
Given a population of solutions, our model is designed to generate new solutions that potentially improve function values. The dual generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ implement the generative loss $\mathcal{L}_{\text{gen}}$ (Eq. 2 and Eq. 7), while the surrogate model $\tilde{f}$ addresses the optimization loss $\mathcal{L}_{\text{opt}}$. Initially, the surrogate model is established and remains constant throughout the training process; $\bm{\gamma}$ processes inferior solutions to generate estimated superior ones, and $\bm{\gamma}^{\prime}$ operates in reverse, both facilitating the computation of the similarity loss $\mathcal{L}_{\text{sim}}$. Subsequently, the solutions generated by $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are reconstructed by cycling back to each other, facilitating the computation of reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5). Additionally, the generated solutions ($\bm{\gamma}(\bm{p})$ and $\bm{\gamma}^{\prime}(\bm{q})$) may also contribute to the optimization loss $\mathcal{L}_{\text{opt}}$ when evaluated against the surrogate. This configuration facilitates the transfer of optimization landscape estimation from the surrogate to the generative models via back-propagation, enabling $\bm{\gamma}$ to autonomously generate optimized populations.
In summary, the architecture integrates a surrogate model $\tilde{f}$, alongside a generative model $\bm{\gamma}$ and its pseudo-inverse $\bm{\gamma}^{\prime}$. $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are specifically trained for generating and reconstructing the superior and inferior solutions, respectively, while $\tilde{f}$ focuses on the estimation of the optimization landscape. Together, their synergistic training significantly enhances $\bm{\gamma}$’s ability to generate optimized solutions.
@@
The generative loss $\mathcal{L}_{\text{gen}}$ comprises two components: a similarity loss $\mathcal{L}_{\text{sim}}$, which encourages consistent latent representations between paired samples, and a reconstruction loss $\mathcal{L}_{\text{rec}}$, which enforces structural consistency in the generated outputs:
@@
The optimization loss $\mathcal{L}_{\text{opt}}$ consists of an expectation loss $\mathcal{L}_{\text{mean}}$, which penalizes poor average fitness among generated samples, and a variance loss $\mathcal{L}_{\text{std}}$, which promotes diversity:
@@
In generative learning, the similarity loss $\mathcal{L}_{\text{sim}}$ is widely employed to ensure that generated outputs closely resemble target samples [76, 75, 77, 78]. For our generative model $\bm{\gamma}$, which transforms a source solution $\bm{p}$ to a target solution $\bm{q}$, we adopt a mean squared loss:
@@
This aligns with conventional generative paradigms, where models learn mappings from latent or random variables to target domains [74, 79, 80]. However, when applied to optimization problems, such models often struggle with the intrinsic trade-off between exploration and exploitation, especially within the EA framework [4, 81, 82].
Alternatively, domain adaptation models, which transform solutions between source and target domains [83, 84, 75, 85], have shown promising results in certain optimization scenarios [82, 86]. Nonetheless, without reconstruction constraints, these models are susceptible to mode collapse [76, 77, 78], compromising training stability. To mitigate this issue, we introduce a paired generative model $\bm{\gamma}^{\prime}$ that serves as a pseudo-inverse of $\bm{\gamma}$. This yields a reconstruction loss:
@@
The generative loss $\mathcal{L}_{\text{gen}}$, particularly the reconstruction term, plays a pivotal role in stabilizing training and preserving semantic fidelity. This is crucial in optimization contexts, where the model must accurately learn transformations from inferior to superior regions of the search space without drifting off-manifold. To ensure this, the weight $\lambda_{1}$ associated with the reconstruction loss is set to dominate the loss function, emphasizing faithful reconstruction over mere proximity.
@@
This formulation adheres to the principles of domain-adaptive generative learning, where $\bm{p}$ and $\bm{q}$ are samples from source (inferior) and target (superior) domains, respectively. The dataset construction process is detailed in Section III-B, and an alternative interpretation of Eq. 7 from a probabilistic (Gaussian) perspective is provided in Supplementary Document A.
Incorporating our generative model into the EA framework requires transforming traditional optimization objectives into suitable loss function terms for complete data-driven process. To some extent, this is similar to the infill criteria in the Surrogate-Assisted Evolutionary Algorithms (SAEAs), where the surrogate model is used to select the most promising solution(s) from a population of candidates [68]. Such strategies have been widely studied in the literature, and one of the most common approach is to use the Upper/Lower Confidence Bound (UCB/LCB) [41] as the infill criteria [65]:
@@
where $\alpha$ is the UCB/LCB strategy to be applied to each candidate $\bm{p}$, $\tilde{f}$ is the surrogate of the objective function which we are trying to minimize and $\lambda_{2}$ is a hyperparameter as mentioned in Eq. 3. In traditional SAEAs, all candidate solutions are evaluated against the Eq. 8 to find the one(s) with the minimal value(s) of $\alpha(\bm{p})$, which is then evaluated by the true objective function.
@@
However, the conventional use of acquisition functions such as UCB or LCB is not directly applicable to our generative model $\bm{\gamma}$. This is because the acquisition strategy must be adapted to incorporate both the surrogate model and the generative model. Inspired by principles from meta-learning [87, 88], we propose to reinterpret the acquisition criterion as a training loss for the generative model. Without loss of generality, we define the generative model as $\bm{\gamma}(\cdot;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, parameterized by $\bm{\phi}$, which generates a new solution $\bm{x}_{t+1}$ from an existing input $\bm{x}$:
@@
In this formulation, the acquisition function $\alpha(\cdot)$ is used to define the optimization objective for training the parameters $\bm{\phi}$ of the generative model. Through this transformation, the task of maximizing the acquisition function is seamlessly integrated into the learning process of the generative model, allowing it to internalize exploration-exploitation trade-offs during training.
Thus, by replacing the objective function $f$ with it surrogate model $\tilde{f}$, we derive the optimization loss of UCB/LCB as:
@@
where $\mathbb{E}$ and $\mathbb{D}$ represent the mean and variance of the surrogate model’s output, respectively. This loss ties the generative model’s parameters directly to the surrogate’s predictions, facilitating a robust integration of surrogate model into a generative modeling context. However, since the transformation of the infill criteria to the optimization loss is not guaranteed to be lossless, the inherent balance between exploration and exploitation may be compromised. Hence, we introduce an explicit variance optimization loss composed of UCB/LCB and Knowledge Gradient (KG) [89]:
@@
where $\mathbb{C}$ is the correlation function. Since the loss function is to be minimized, the existence of correlation function requests that the parameter $\lambda_{2}$ is negative, which is not necessarily the case in Eq. 10. Supplementary Document A will elaborate the details of the justification of Eq. 11.
It is important to note that, unlike in conventional SAEAs, the surrogate model $\tilde{f}$ in our framework is not directly used to evaluate the quality of generated solutions. Instead, it serves as a guidance mechanism during the training of the generative model. Furthermore, given the wide variety of infill criteria proposed in the literature, the acquisition function in Eq. 8 can be flexibly substituted with alternative strategies. This substitution naturally leads to different formulations of the optimization loss $\mathcal{L}_{\text{opt}}$, thereby allowing our framework to accommodate various exploration-exploitation trade-offs.
@@
The training procedure is summarized in Algorithm 2. It begins with training the surrogate model $\tilde{f}$ on the elite dataset $\mathcal{D}_{\text{SM}}$, after which the surrogate is frozen to ensure stable guidance throughout the optimization process. The paired generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are subsequently optimized using mini-batches sampled from the paired dataset $\mathcal{D}_{\text{GM}}$, which is constructed during the data preparation phase (Algorithm 1).
@@ III-D Population Generation
Similar to the reproduction mechanism in traditional EAs, the population generation phase in EvoGO aims to create new candidate solutions based on the current population. However, instead of relying on handcrafted heuristic operators, EvoGO utilizes the learned generative model $\bm{\gamma}$ for solution generation.
@@
5:  $\bm{x}_{i}^{t+1}\leftarrow\bm{\gamma}(\bm{x}_{i}^{t};\bm{\phi}^{*})$ // “$\bm{\gamma}$” for trained generative model
@@
As illustrated in Algorithm 3, this phase employs the trained generative model $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ to transform inferior solutions into superior ones, thereby generating new candidate solutions from the current population $\mathcal{D}_{t}$. Each generated solution $\bm{x}^{t+1}_{i}$ is evaluated using the true objective function $f$ to obtain its fitness value $y^{t+1}_{i}$, forming the next-generation population $\mathcal{D}_{t+1}$.
Due to the inherently parallelizable nature of the generative model, this step can be fully parallelized, provided that the objective function $f$ supports parallel evaluation—thus significantly improving computational efficiency.

### v2

@@ III-A Overall Framework
1:Input: historical data $\mathcal{D}$, dataset split factor $\eta\in(0,1)$, sliding window factor $\epsilon>=0$, data augmentation threshold NN, and factor $\rho$
@@
6: /******** Data Augmentation (Optional) ********/
@@
As illustrated in Fig. 1, EvoGO comprises three key, sequentially executed phases: data preparation, model training, and population generation. To streamline the evolutionary process and eliminate reliance on complex handcrafted rules, EvoGO adopts an end-to-end generative paradigm, replacing the heuristic components commonly used in traditional GMEAs.
This generative architecture directly addresses two core challenges inherent in existing GMEAs: data scarcity and objective misalignment. Specifically, the data preparation phase incorporates a data augmentation strategy to mitigate the risk of overfitting caused by limited population data, thereby enhancing training diversity without introducing additional evaluation costs. Furthermore, in the model training phase, a composite model architecture and a tailored loss function are employed to align the generative objectives with the goals of evolutionary optimization, thereby resolving the objective misalignment issue. The time and space complexity of EvoGO are analyzed in detail in Supplementary Document A.
@@ III-B Data Preparation
The data preparation phase is pivotal for constructing high-quality and well-structured training data that is sufficient for both the surrogate model and the generative model, particularly under limited evaluation budgets. Achieving high generalization capabilities for the generative model requires rich, diverse input data, whereas the surrogate model demands highly promising samples for accurate fitness estimation. Consequently, this phase is designed to manage the trade-off between diversity and fidelity to maximize data utility.
@@
  • Elite Selection: A high-quality dataset $\mathcal{D}_{\text{SM}}$ is constructed by selecting individuals with the highest fitness values from all historical populations $\mathcal{D}_{0}\cup\cdots\cup\mathcal{D}_{t}$. This process ensures accurate surrogate modeling by focusing on representative and promising solutions.
  • Data Augmentation (Optional): When the evaluation budget is limited, requiring the algorithm to operate with a small population, a learning-based data augmentation strategy is adopted. Specifically, a Variational AutoEncoder (VAE) is used to generate diverse synthetic samples that resemble the current population. This process enriches the generative model dataset $\mathcal{D}_{\text{GM}}$, thereby alleviating the data scarcity issue under restricted evaluations.
  • Paired Set Construction: Instead of enlarging the dataset further, this step enhances data utilization by constructing informative sample pairs. The surrogate model estimates the fitness of all candidates, after which the samples are ranked and divided into superior ($\mathcal{D}_{\text{GM}^{+}}$) and inferior ($\mathcal{D}_{\text{GM}^{-}}$) subsets according to a predefined ratio $\eta$. The Cartesian product of the two subsets forms the paired training set $\mathcal{D}_{\text{GM}}$, enabling the generative model to learn both effective descent directions and adaptive step magnitudes through relational supervision.
@@ III-C Model Training
The model training phase focuses on establishing a composite model architecture integrated with a dedicated composite loss function. This architecture enables directional learning and optimization-aware training through an online process, where lightweight training is conducted during each evolutionary generation. Although the composite structure involves three model components, only the forward generative model $\bm{\gamma}$ is ultimately used during the population generation stage. The subsequent text details the overall training procedure, beginning with the model architecture and followed by the formal loss formulation.
@@
Inspired by Generative Adversarial Networks (GANs) [79] and CycleGANs [80], we develop a composite-model-based GMEA framework (see Fig. 2) that integrates a surrogate model $\tilde{f}$ with two complementary generative components. These components are a forward generative model $\bm{\gamma}$, which maps inferior solutions to superior ones, and a pseudo-inverse model $\bm{\gamma}^{\prime}$, which performs the reverse mapping. Unlike conventional GAN-like architectures that rely on discriminators, the proposed framework employs the surrogate model to enforce optimization-aware learning, thereby aligning generative behaviors with the underlying objective landscape.
Given an evolving population, the composite architecture is designed to generate new solutions that possess improved function values. The dual generative models, $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$, implement the generative loss $\mathcal{L}_{\text{gen}}$ (Eq. 2 and Eq. 7), whereas the surrogate model $\tilde{f}$ addresses the optimization loss $\mathcal{L}_{\text{opt}}$. Specifically, $\bm{\gamma}$ processes inferior solutions to generate estimated superior ones, and $\bm{\gamma}^{\prime}$ operates in reverse, both facilitating the computation of the similarity loss $\mathcal{L}_{\text{sim}}$.
The integrated architecture further facilitates the transfer of optimization landscape estimation from the surrogate model to the dynamic generative components. Initially, the surrogate model $\tilde{f}$ is established and remains constant throughout the training process. Subsequently, the solutions generated by $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are reconstructed by cycling back to each other, which facilitates the computation of the reconstruction loss $\mathcal{L}_{\text{rec}}$ (Eq. 5). Crucially, the generated solutions ($\bm{\gamma}(\bm{p})$ and $\bm{\gamma}^{\prime}(\bm{q})$) are also evaluated against the surrogate, contributing to the optimization loss $\mathcal{L}_{\text{opt}}$. This configuration allows for the transfer of optimization landscape information via back-propagation, thereby enabling $\bm{\gamma}$ to autonomously generate optimized populations.
In summary, the architecture synergistically integrates a static surrogate model $\tilde{f}$ with a dynamic generative model $\bm{\gamma}$ and its pseudo-inverse $\bm{\gamma}^{\prime}$. $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are specifically trained for generating and reconstructing the superior and inferior solutions, respectively. Conversely, $\tilde{f}$ focuses exclusively on the estimation of the optimization landscape. This synergistic training substantially enhances the practical ability of $\bm{\gamma}$ to generate optimized solutions.
@@
The generative loss $\mathcal{L}_{\text{gen}}$ comprises two core components. The first is the similarity loss $\mathcal{L}_{\text{sim}}$, which encourages consistent latent representations between paired samples. The second is the reconstruction loss $\mathcal{L}_{\text{rec}}$, which enforces structural consistency in the generated outputs:
@@
The overall optimization loss $\mathcal{L}_{\text{opt}}$ is formulated as the combination of two components: the expectation loss $\mathcal{L}_{\text{mean}}$, which penalizes poor average fitness among generated samples, and the variance loss $\mathcal{L}_{\text{std}}$, which promotes diversity:
@@
The similarity loss $\mathcal{L}_{\text{sim}}$ is conventionally employed in generative learning paradigms to ensure that generated outputs closely resemble the target samples [81, 80, 82, 83]. Specifically, for the generative model $\bm{\gamma}$, which transforms a source solution $\bm{p}$ to a target solution $\bm{q}$, we adopt a mean squared loss:
@@
This aligns with conventional generative paradigms, where models learn mappings from latent or random variables to target domains [79, 84, 85]. However, when applied to optimization problems, such models often struggle with the intrinsic trade-off between exploration and exploitation, especially within the EA framework [5, 86, 87]. Alternatively, domain adaptation models, which transform solutions between source and target domains [88, 89, 80, 90], have shown promising results in certain optimization scenarios [87, 91]. Nonetheless, without reconstruction constraints, these models are susceptible to mode collapse [81, 82, 83], thereby compromising training stability. To mitigate this issue, we introduce a paired generative model $\bm{\gamma}^{\prime}$ that serves as a pseudo-inverse of $\bm{\gamma}$. This yields a reconstruction loss:
@@
The generative loss $\mathcal{L}_{\text{gen}}$, particularly the reconstruction term, plays an essential role in stabilizing training and preserving semantic fidelity. This mechanism is crucial in optimization contexts, where the model must accurately learn transformations from inferior to superior regions of the search space without drifting off-manifold. To ensure this, the weight $\lambda_{1}$ associated with the reconstruction loss is set to dominate the loss function, emphasizing faithful reconstruction over mere proximity.
@@
This formulation adheres to the principles of domain-adaptive generative learning, where $\bm{p}$ and $\bm{q}$ are samples from source (inferior) and target (superior) domains, respectively. The dataset construction process is detailed in Section III-B, and an alternative interpretation of Eq. 7 from a probabilistic perspective is provided in Supplementary Document B.
Integrating the generative model into the EA framework necessitates transforming traditional optimization objectives into suitable loss function terms for a complete data-driven process. This requirement is analogous to the infill criteria used in SAEAs, where the surrogate model selects the most promising solutions from a population of candidates [73]. Such selection strategies have been widely studied in the literature. One of the most common approaches is to use the Upper/Lower Confidence Bound (UCB/LCB) [43] as the infill criteria [68]:
@@
Traditional Surrogate-Assisted Evolutionary Algorithms (SAEAs) utilize the infill criteria, $\alpha(\bm{p})$, to guide the selection of candidate solutions $\bm{p}$ for high-fidelity evaluation. Specifically, $\tilde{f}$ denotes the surrogate of the objective function being minimized, and $\lambda_{2}$ is a hyperparameter defined in Eq. 3. All candidate solutions are evaluated using the criterion established in Eq. 8 to identify those with minimal $\alpha(\bm{p})$ values, which are subsequently passed to the true objective function for final evaluation.
@@
However, the conventional application of infill criteria, such as UCB or LCB, is not directly applicable to our proposed generative model $\bm{\gamma}$. This is because the standard infill criteria must be adapted to cohesively incorporate both the surrogate model and the generative model simultaneously. Inspired by meta-learning principles [92, 93], we propose to reinterpret the infill criteria as an optimization loss for training the generative model. Without loss of generality, we define the generative model as $\bm{\gamma}(\cdot;\bm{\phi}):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, parameterized by $\bm{\phi}$, which generates a new solution $\bm{x}_{t+1}$ from an existing input $\bm{x}$:
@@
In this formulation, the infill criteria $\alpha(\cdot)$ explicitly defines the optimization objective used for training the parameters $\bm{\phi}$ of the generative model. Through this transformation, the task of minimizing the infill criteria is seamlessly integrated into the learning process of $\bm{\gamma}$, thereby allowing the model to internalize the required exploration-exploitation trade-offs during training.
Thus, by replacing the underlying objective function $f$ with its surrogate model $\tilde{f}$ in this generalized framework, we derive the resulting optimization loss for UCB/LCB as:
@@
where $\mathbb{E}$ and $\mathbb{D}$ represent the mean and variance of the surrogate model’s output, respectively. This loss ties the generative model’s parameters directly to the surrogate’s predictions, facilitating a robust integration of the surrogate model into a generative modeling context. However, since the transformation of the infill criteria to the optimization loss is not guaranteed to be lossless, the inherent balance between exploration and exploitation may be compromised. Hence, we introduce an explicit variance optimization loss composed of UCB/LCB and Knowledge Gradient (KG) [94]:
@@
where $\mathbb{C}$ is the correlation function. Since the loss function is to be minimized, the existence of the correlation function requests that the parameter $\lambda_{2}$ is negative, which is not necessarily the case in Eq. 10. Supplementary Document B will elaborate the details of the justification of Eq. 11.
Unlike in conventional SAEAs, the surrogate model $\tilde{f}$ in our framework is not directly used to evaluate the quality of generated solutions. Instead, it serves only as a guidance mechanism during the training of the generative model. Furthermore, given the wide variety of infill criteria proposed in the literature, the infill criteria defined in Eq. 8 can be flexibly substituted with alternative strategies. This substitution naturally leads to different formulations of the optimization loss $\mathcal{L}_{\text{opt}}$, thereby allowing our framework to accommodate various exploration-exploitation trade-offs.
@@
The training procedure is summarized in Algorithm 2. It begins with training the surrogate model $\tilde{f}$ on the elite dataset $\mathcal{D}_{\text{SM}}$, after which the surrogate is frozen to ensure stable guidance throughout the optimization process. Subsequently, the paired generative models $\bm{\gamma}$ and $\bm{\gamma}^{\prime}$ are optimized using mini-batches sampled from the paired dataset $\mathcal{D}_{\text{GM}}$, which is constructed during the data preparation phase (Algorithm 1).
@@ III-D Population Generation
Similar to the reproduction mechanism employed in traditional EAs, the population generation phase in EvoGO aims to create new candidate solutions based on the current population. However, instead of relying on handcrafted heuristic operators, EvoGO utilizes the learned generative model $\bm{\gamma}$ for solution generation.
@@
5:  $\bm{x}_{i}^{t+1}\leftarrow\bm{\gamma}(\bm{x}_{i}^{t};\bm{\phi}^{*})$ // $\bm{\gamma}$” for trained generative model
@@
As illustrated in Algorithm 3, this phase employs the trained generative model $\bm{\gamma}(\cdot;\bm{\phi}^{*})$ to transform inferior solutions into superior ones, thereby generating new candidate solutions from the current population $\mathcal{D}_{t}$. Each generated solution $\bm{x}^{t+1}_{i}$ is subsequently evaluated using the true objective function $f$ to obtain its fitness value $y^{t+1}_{i}$, forming the next-generation population $\mathcal{D}_{t+1}$. Due to the inherently parallelizable nature of both the generative model and the objective function $f$, this entire population generation and evaluation step can be fully parallelized, which substantially improves computational efficiency.